OpenStudy (anonymous):

Solve the following equation: x-4=6

OpenStudy (anonymous):

You add 4 to both sides and get x=10. All you have to do is isolate x in this equation.

OpenStudy (anonymous):

you can add a constant to both sides of an equation, add the constant four to both sides and simplify

OpenStudy (anonymous):

\[x-4=6\] \[x=6+4 \] \[x=10\]

OpenStudy (anonymous):

X-4=6 X=6+4 X=10